what is the surface area of a marble slab that is measured 5 meters high, by 7 meters wide, and 2 meters deep?
Answer by ankor@dixie-net.com(12692)

what is the surface area of a marble slab that is measured 5 meters high, by 7 meters wide, and 2 meters deep?
:
2(7*2) + 2(7*5) + 2(5*2) =
28 + 70 + 20 = 118 sq meters